OpenStudy (anonymous):

I've been stuck on this question for an hour now. Can you please help me do it and understand it? The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

So far I have: \[A = L*W\] and that'll be \[280 = (l -5) (w + 1)\].

OpenStudy (campbell_st):

so in part A 280 = l*w so l= 280/w then using the new dimensions you have posted and making a substitution l - 5 = w + 1 since the sides of the square are equal make the substitution \[\frac{280}{w} - 5 = w + 1\] multiply every term by w \[280 - 5w = w^2 + w\] or \[w^2 + 6w - 280 = 0\] just solve the quadratic for w. you will have 2 answers, just use the positive... as a measurement can't be negative hope it helps

OpenStudy (campbell_st):

the quadratic equation can be factored.

OpenStudy (campbell_st):

or use the general quadratic formula to get the solutions

OpenStudy (anonymous):

Thank you! I found that the width is 14. Am I correct?

OpenStudy (campbell_st):

that's correct...

OpenStudy (anonymous):

Thank you again! :)